# Your Language Model Can Secretly Write Like Humans: Contrastive Paraphrase Attacks on LLM-Generated Text Detectors

### v2

@@ Appendix C Additional Results
No Attack R1-Distill-32B QwQ-32B Qwen2.5-7B GLM-4-9b-hf Qwen2.5-72B
XSum Fast-DetectGPT 95.33 24.33 11.33 17.33 31.33 17.00
TOCSIN 98.00 42.67 10.00 26.00 87.00 26.67
R-Detect 49.33 6.33 00.00 1.33 38.33 0.00
SQuAD Fast-DetectGPT 93.50 26.83 9.50 33.33 31.33 27.50
TOCSIN 92.67 42.00 17.33 26.00 81.33 25.33
R-Detect 58.00 13.00 0.00 1.67 39.67 0.00
LongQA Fast-DetectGPT 86.00 14.67 11.33 14.00 27.33 11.33
TOCSIN 88.67 26.00 2.00 18.00 78.00 16.00
R-Detect 74.33 22.67 0.00 3.00 47.33 3.00
@@
Results of More LLMs as paraphrasers. To validate the universality of the proposed attack, we next consider more LLMs as the paraphraser, including various model scales and recently prevalent reasoning-based models. Specifically, we consider Deepseek R1-Distill-32B DeepSeek-AI (2025), QwQ-32B Team (2025), Qwen2.5-7B Qwen (2024), and GLM-4-9B-hf GLM (2024). The quantitative results in Table 5 show the effectiveness of the proposed CoPA across various LLMs. Note that the QwQ-32B generally achieves the best performance. However, the reasoning-based models require significantly more inference time than regular models. Hence, we choose the Qwen2.5-72B to balance effectiveness and efficiency.
@@ Appendix E Details about Machine Prompt
The prompts used in Figure 3 are as follows:
@@
[…]
@@
Appendix F An Adaptive Defense Strategy
To alleviate the proposed threat, we implement an adaptive defense that adversarially trains OpenAI’s LLM-text classifier RoBERTa-large. We fine-tune the model for 10 epochs using 5k human texts and 5k machine texts (including both the original machine texts and those paraphrased by CoPA). We present the optimal performance at a proportion of 50% CoPA-paraphrased samples within the 5k machine texts in Table 7.
Appendix G Evaluation on Text Quality
Apart from the attack effect, it is necessary to analyze the text quality of paraphrased sentences. Specifically, we provide a comprehensive evaluation of natural fluency and semanticconsistency with additional key metrics, including text perplexity, GPT4-assisted evaluation, and human study. To conduct the human evaluation, we choose GPT-3.5-turbo as the source model and randomly select 100 pairs of texts from each dataset for human annotators. The evaluation criteria generally align with those in Dipper Krishna et al. (2023), where we recruit 10 native English speakers from Amazon Mechanical Turk (MTurk) to perform the evaluation. We report the average scores to reduce subjective biases in Table 8. The results indicate that the proposed method produces paraphrased texts with lower perplexity than authentic human-written texts, while achieving substantially better fluency and semantic consistency compared to those generated by Dipper, in both terms of GPT-4 assited and human evaluation.

### v3

@@ Appendix C Additional Results
No Attack R1-Distill-32B QwQ-32B GLM-4-9b-hf Qwen2.5-72B
XSum Fast-DetectGPT 95.33 24.33 11.33 31.33 17.00
TOCSIN 98.00 42.67 10.00 87.00 26.67
R-Detect 49.33 6.33 00.00 38.33 0.00
SQuAD Fast-DetectGPT 93.50 26.83 9.50 31.33 27.50
TOCSIN 92.67 42.00 17.33 81.33 25.33
R-Detect 58.00 13.00 0.00 39.67 0.00
LongQA Fast-DetectGPT 86.00 14.67 11.33 27.33 11.33
TOCSIN 88.67 26.00 2.00 78.00 16.00
R-Detect 74.33 22.67 0.00 47.33 3.00
@@
Results of More LLMs as paraphrasers. To validate the universality of the proposed attack, we next consider more LLMs as the paraphraser, including various model scales and recently prevalent reasoning-based models. Specifically, we consider Deepseek R1-Distill-32B DeepSeek-AI (2025), QwQ-32B Team (2025), and GLM-4-9B-hf GLM (2024). The quantitative results in Table˜5 show the effectiveness of the proposed CoPA across various LLMs. Note that the QwQ-32B generally achieves the best performance. However, the reasoning-based models require significantly more inference time than regular models. We choose the Qwen2.5-72B to balance effectiveness and efficiency.
@@
Ablation study of the adaptive truncation mechanism. To avoid penalizing the probabilities of reasonable and valid tokens, we incorporate an adaptive truncation mechanism to constrain the output token distribution. We provide an ablation study to investigate its influence. As observed in Table 7, by truncating the token distribution within a reliable token candidate pool, we can improve the text quality of the generated sentences while maintaining attack effectiveness.
Table 7: Ablation on the adaptive truncation technique.
Method Sim Fast-DetectGPT TOCSIN R-Detect
w/o truncation 93.33 19.67 31.33 5.33
CoPA 94.00 17.00 26.67 4.67
@@ Appendix E Details about Machine Prompt
The prompts used in Figure˜3 are as follows:
@@
Appendix F An Adaptive Defense Strategy
To alleviate the proposed threat, we implement an adaptive defense that adversarially trains OpenAI’s LLM-text classifier RoBERTa-large. We fine-tune the model for 10 epochs using 5k human texts and 5k machine texts (including both the original machine texts and those paraphrased by CoPA). We present the optimal performance at a proportion of 50% CoPA-paraphrased samples within the 5k machine texts in Table˜8.
@@ Appendix F An Adaptive Defense Strategy
Appendix G Human Prompts for Different Contexts
To better handle the different demands of contexts, we conduct a preliminary study by designing two alternative prompts as follows:
Human Prompt I for General Scenarios Rewrite the following INPUT in human style, with varying sentence structures and replace common terms with nuanced synonyms. Maintain the meaning and avoid any repetition.
Human Prompt II for Academic Context Rewrite the following INPUT in a human-written academic style, with varying sentence structures and replace common terms with nuanced synonyms. Maintain the meaning and avoid any repetition.
Here are the corresponding attack results:
[…]
Prompt Fast-DetectGPT TOCSIN RoBERTa R-Detect
Prompt I 24.67 43.33 15.33 18.67
Prompt II 24.33 39.67 5.33 14.33
CoPA 17.00 26.67 22.67 4.67
The quantitative results in Tab. 1 and demonstrations of paraphrased sentences show that CoPA is able to flexibly combine with prompts of various styles while maintaining high effectiveness. Users may design their own prompts based on our provided template to adapt to diverse writing styles.
Appendix H Evaluation on Text Quality
Apart from the attack effect, it is necessary to analyze the text quality of paraphrased sentences. Specifically, we first conduct a deeper linguistic analysis of the output texts in Table 10, including TTR (Type-Token Ratio) and MTLD (Measure of Textual Lexical Diversity) metrics for lexical diversity, Div_\_syn Guo et al. (2024b) for syntactic variability. The results demonstrate the high quality of our generated sentences.
Table 10: Analysis of paraphrased sentences from different methods based on the XSum dataset.
Method TTR↑ MTLD↑ Div_\_syn↑
No Attack 0.5734 111.4913 0.4700
Dipper 0.5511 74.8257 0.4648
CoPA 0.6746 145.8134 0.5593
@@ Appendix H Evaluation on Text Quality
$uparrow Natural fluency\uparrow Consistency\uparrow$
@@
We also provide a comprehensive evaluation of natural fluency and semantic consistency with additional key metrics, including text perplexity, GPT4-assisted evaluation, and human study. To conduct the human evaluation, we choose GPT-3.5-turbo as the source model and randomly select 100 pairs of texts from each dataset for human annotators. The evaluation criteria generally align with those in Dipper Krishna et al. (2023), where we recruit 10 native English speakers from Amazon Mechanical Turk (MTurk) to perform the evaluation. We report the average scores to reduce subjective biases in Table˜11. The results indicate that CoPA produces paraphrased texts with lower perplexity than authentic human-written texts, while achieving substantially better fluency and semantic consistency compared to those generated by Dipper, in both terms of GPT-4 assited and human evaluation.